A part-time instructor is receiving $950 per
credit taught. If the instructor receives a 4% increase,
how much will the new per credit compensation be

Answer by Maths68(1474)

New Per credit compensation = Old compensation + increase%
New Per credit compensation = 950+(950)(4/100)
New Per credit compensation = 950+(950)(1/25)
New Per credit compensation = 950+38
New Per credit compensation = 988